How far does a sled travel in 3 hours while accelerating from 10 miles/hour to 14 miles/ hour?


1
Expert's answer
2021-11-10T10:09:04-0500
"d=\dfrac{1}{2}(v_0+v)t,""d=\dfrac{1}{2}\times(10\ \dfrac{mi}{h}+14\ \dfrac{mi}{h})\times3\ h=36\ mi."